Supplementary angles are two angles whose measures have the sum of 180 degrees. Angles X and Y are supplementary, and the measure of angle X is 24 degrees greater than the measure of angle Y. Find the measures of angles X and Y.
Answer by edjones(7311)

x=y+24
x+y=180
.
y+24+y=180 Substitute y+24 for x.
2y=156
y=78 degrees
x=102 degrees
.
Ed